OpenStudy (anonymous):

If AC = 15 centimeters, and BC = 12 centimeters, which does AB equal? A.6 centimeters B.8 centimeters C.9 centimeters D.10 centimeters

OpenStudy (anonymous):

@iGreen

OpenStudy (k_lynn):

Do you have a line that looks something like this? http://prntscr.com/6zwr2l

OpenStudy (k_lynn):

@david812

OpenStudy (anonymous):

yes i do im sorry i wasnt on @k_lynn

OpenStudy (hockeychick23):

is this for legs of a triangle? Because if it is then you could use the Pythagorean Theorem which is a^2+b^2=c^2 so you just fill in the numbers: 12^2+b^2=15^2 --> 144+b^2=225. Then subtract 144 from both sides: b^2+144-144=225-144 --> b^2=81. Then you take the square root of both sides and b=9

OpenStudy (k_lynn):

ok @david812 http://prntscr.com/709f5z What do you think you should do? Let me know if my picture is somewhat accurate.